OpenStudy (anonymous):

Solve the logarthmic equation. 2 log3(x + 4) - log39 = 2

OpenStudy (anonymous):

does it help to know that \(\log_3(9)=2\) ?

OpenStudy (anonymous):

that is because \(3^2=9\) so you have \[2\log_3(x+4)-2=2\] or \[2\log_3(x+4)=4\] or \[\log_3(x+4)=2\]

OpenStudy (anonymous):

that means \(3^2=x+4\) and that is easy to solve

OpenStudy (anonymous):

\[\log_{3} \left( x+4 \right)^2-\log_{3} 9=2\] \[\log_{3} \frac{ \left( x+4 \right)^2 }{ 9 }=2\] \[\log_{3} \left( \frac{ x+4 }{3 } \right)^2=2,2\log_{3} \frac{ x+4 }{ 3 }=2\] \[\log_{3} \frac{ x+4 }{3 }=\frac{ 2 }{ 2 }=1,\frac{ x+4 }{3 }=3^1=3,x+4=9,x=?\]